The length of a rectangle can be found using the following formula.



What is the length of a rectangle that has a width of 2.5 centimeters and an area of 20.5 square centimeters?
1
Expert's answer
2020-11-09T19:46:19-0500

Area of rectangle equal the product of its width and length.

S=wxl, where

w - the width of rectangle.

l - the length of rectangle.

S - the area of rectangle.

So, l=S/w =20.5/2.5=8.2

Answer: The length of rectangle is 8.2 centimetres.